# Exact Work Distribution and Jarzynski’s Equality of a Relativistic Particle in an Expanding Piston

### v1

@@ III.5 Limit of Fast-moving Piston
$\begin{split}\left\langle e^{\beta W}\right\rangle&=P_{0}\cdot e^{\beta\cdot 0}\\ &+\int_{v_{p}}^{1}\mathrm{d}v\frac{\varphi_{1}(v)e^{-\frac{\beta}{\sqrt{1-v^{2}}}}}{2K_{1}(\beta)(1-v^{2})^{\frac{3}{2}}}\cdot e^{\beta W}.\end{split}$ (42)
The first term is approximately equal to 1 ($P(W>0)\ll 1$, $P_{0}\approx 1$). From the value of $\exp{(-\beta\Delta F)}=1+v_{p}\tau$, one can expect that, although the probability $P(W>0)$ is vanishingly small, the contribution of the second term is non-zero. This result can be demonstrated transparently in the low-temperature limit. Noticing that
$\begin{split}e^{-\beta(1-v^{2})^{-\frac{1}{2}}}\cdot e^{\beta W}=&e^{-\beta H_{\lambda(0)}}\cdot e^{\beta(H_{\lambda(0)}-H_{\lambda(\tau)})}\\ =&e^{-\beta H_{\lambda(\tau)}}\\ =&e^{-\beta(1-v_{1}^{2})^{-\frac{1}{2}}},\end{split}$ (43)
@@
When $\beta\to\infty$, $e^{-\beta(1-v^{2}_{1}(v))^{-\frac{1}{2}}}$ becomes a Gaussian peak, which is a $\delta$ distribution around $v=2v_{p}/(1+v_{p}^{2})$,
$\begin{split}e^{-\beta(1-v^{2}_{1}(v))^{-\frac{1}{2}}}\approx&e^{-\beta}\sqrt{\frac{2\pi}{\beta}}\delta(v_{1}(v))\\ =&e^{-\beta}\sqrt{\frac{2\pi}{\beta}}\left[1-\left(\frac{2v_{p}}{1+v_{p}^{2}}\right)^{2}\right]\delta\left(v-\frac{2v_{p}}{1+v_{p}^{2}}\right).\end{split}$ (44)
Together with Eq. (38) and Eq. (28), the second term in Eq. (42) becomes
@@
This result demonstrates that when the piston moves at a very large $v_{p}$, particles with high initial velocities around $2v_{p}/(1+v_{p}^{2})$ contribute most significantly to the exponential work even if the probability is extremely small. Please note that in Newtonian mechanics particles with initial velocities around $2v_{p}$ contribute most significantly to the exponential work [4][10], even if the probability is extremely small. It can be seen that the results of the relativistic piston model recover those of the non-relativistic piston model as expected.
IV Discussion and Summary
@@
Let us look further into the main results we obtained. We see that, as a consequence of the relativistic energy-velocity relation, the trapezoid-shaped work distribution no longer has a series of zeros. Moreover, the number of peaks becomes finite because the speed of light places an upper bound on all speeds. The apparent paradox of the fast-moving piston in Ref. [3] can be reformulated as when the piston is moving at the speed of light instead of infinity. No particle would be able to catch up with the piston and the average exponential work is unity whereas the free energy change is non-zero. We would like to point out that although light-like worldlines exist, we can not make it stop before and after the moving time period, because such a worldline configuration would violate causality. The best we can do is to take the limiting process of letting the speed of the piston approach the speed of light. Then the order of limit becomes crucial, as we have to integrate out the work $W$ to infinity, and then take the speed limit [4]. This limiting procedure ensures the validity of the JE.
@@
In order to observe a distinct deviation of the relativistic work distribution from its non-relativistic limit, the speed of the piston should be large enough and the temperature of the ideal gas should be extremely high. Take hydrogen atoms as an example, to observe the features of the relativistic work distribution, the piston should be as fast as about $3\times 10^{6}\mathrm{m/s}$, and the temperature of the atoms should be about $10^{12}\mathrm{K}$. When the speed of the piston is about $3\times 10^{5}\mathrm{m/s}$ (faster than the Parker Solar Probe, which is the fastest object human ever built) and the temperature of the atoms is $10^{11}\mathrm{K}$($10^{4}$ times hotter than the central temperature of the Sun), the deviation of the relativistic work distribution from the non-relativistic result becomes barely detectable. Even in such a circumstance, the boundary condition is still difficult to be realized. Because the energy scale of the kinetic energy of the atoms is much larger than the energy scale of the chemical bond, the boundary can not be built by any materials we have already discovered. Based on these facts, we conclude that it is difficult to detect the relativistic effects of the work distribution of ideal gas in a piston system with the current experimental techniques.
@@
In summary, we study a simple model of a piston and ideal gas in the framework of the special theory of relativity. We obtain an analytical result of the work distribution (27) and verify the JE. Using our result it is possible to see the deviation of relativistic work distribution (27) from the non-relativistic one (39) [3]. In principle, these relativistic corrections become non-negligible in the high-temperature and fast-speed limit. However, the range of parameters where relativistic effects are observable would already be far beyond the current experimental techniques. Our results show that the JE holds true in a wide range of systems with generality and serve the pedagogical purpose.
Acknowledgements
@@
This work is supported by the National Natural Science Foundations of China (NSFC) under Grants No. 12375028 and No. 11825501.
Appendix A A Covariant Description of the Trajectory
@@
In Eq. (15), we extend the range of the initial position $x$ from $[0,L]$ to $[-L,L]$ in order to remove the negative sign related to the direction of $v$. We also notice that in Eq. (30), when $t_{n}$ happens to be $\tau$, the critical initial position and the initial velocity satisfy a linear relation. We will introduce a coordinate transformation method to describe the trajectory of a particle, which will provide pictorial intuition of the two facts.
@@
To begin with, let us consider the collision between the particle and the fixed boundary. In the reference frame of the fixed boundary (which is the same as the laboratory frame), the coordinate frame consists of $x$-axis and $t$-axis, and the world line of the moving piston is $l$ (see Fig. 1(a)). We could draw an auxiliary world line of the moving piston, named $l^{\prime}$. The auxiliary world line $l^{\prime}$ and the real world line $l$ are mirror-symmetric about the $t$-axis. The world line of the particle is a polyline line $I-A-B$. Point $I$ denotes the initial condition of the particle, point $A$ denotes the collision event between the particle and the fixed boundary, and point B denotes the next collision event between the particle and the moving piston. Transformation of the particle’s world line works as follows: the auxiliary event of event $B$ is point $B^{\prime}$. $B$ and $B^{\prime}$ are mirror-symmetric about the $t$-axis. The auxiliary world line of the particle is $I-A-B^{\prime}$. The real world line’s $AB$ part and the auxiliary world line’s $AB^{\prime}$ part are also mirror-symmetric about the $t$-axis. As a result of such transformation, the auxiliary world line of the particle $I-A-B^{\prime}$ is a straight line, which eliminates the world line’s direction change during a collision. Meanwhile, in the frame of the fixed boundary, the auxiliary collision event $B^{\prime}$ has the same time coordinate as event $B$. Therefore, we could use an auxiliary event to evaluate the time when a real collision takes place.
@@
We may also map the initial condition $I(x,-v)$ to an auxiliary initial condition $I^{\prime}(-x,v)$, which is the same as the treatment for Eq. (15). The real motion is that a particle starts from the initial position $I$ with a positive initial position $x$ but a negative initial velocity $-v$ (the particle moves away from the moving piston), and then the particle collides with the fixed boundary at event $A$. Meanwhile, the auxiliary motion is that a particle starts from $I^{\prime}$ with a positive initial velocity $v$ but a negative initial position $-x$. The auxiliary particle passes through the fixed boundary at event $A$ without any collision. After event $A$, both the particle and the auxiliary particle move towards the moving piston, and finally collide with the piston at the same event $B$. The rest trajectories are equivalent, for both the particle with the initial condition $(x,-v)$ and the auxiliary particle with the initial condition $(-x,v)$. Because there is no work done during the collision event $A$, it is convenient to extend the range of $x$ from $[0,L]$ to $[-L,L]$ while limiting the range of $v$ from $[-c,c]$ to $[0,c]$ in the calculation of the work distribution.
Figure 2: Transformation of a world line. For details, please see the main text of Appendix A.
@@
After explaining the treatment in Eq. (15), we now continue to the understanding of Eq. (30). Similar to the mirror-symmetric operation in the laboratory reference frame, we could deal with the collision between the particle and the moving piston by carrying out a mirror-symmetric operation in the reference frame of the piston. After a Lorentz transformation back to the laboratory reference frame, the result is shown in Fig. 1(b). Points $I$,$A$,$B$ denote the initial state of the particle, a collision between the particle and the moving piston, and the next collision between the particle and the fixed boundary. $B^{\prime}$ is the auxiliary event of $B$. Just like the case in Fig. 1(a), in the reference frame of the moving piston, the auxiliary world line of the particle is a straight line. Therefore, after a Lorentz transformation the auxiliary world line of the particle is still a straight line $I-A-B^{\prime}$. The auxiliary world line of the fixed boundary is denoted as $t^{\prime}$. The velocity of $t^{\prime}$ is $2v_{p}/(1+v_{p}^{2})$, according to the Lorentz transformation. The intercept of the line $t^{\prime}$ can be determined as follows: the world line of the fixed boundary and moving piston intersect at point $R$ with space and time coordinate $(0,-L/v_{p})$. In the reference frame of the moving piston, the auxiliary world line of the fixed boundary and the moving piston intersect at the same point $R$. Therefore, in the laboratory reference frame, the auxiliary world line of the fixed boundary $t^{\prime}$ also passes through $R$, and the intercept of $t^{\prime}$ must be $-L/v_{p}$.
@@
The mirror-symmetric operation transforms not only the world lines, but also the coordinate frames. In the reference frame of the fixed boundary, the $t$-axis is the world line of the boundary itself. After a coordinate transformation related to a collision with the moving piston, the auxiliary world line of the fixed boundary represents the auxiliary $t^{\prime}$-axis. The auxiliary $x^{\prime}$-axis should be orthogonal to $t^{\prime}$-axis, and the origin of coordinates (point $O$) should be transformed accordingly: $x$-axis and $l$ intersect at $P_{0}$, and thus $x^{\prime}$-axis also passes through $P_{0}$ (see Fig. 2(a). The intersection of the auxiliary $x^{\prime}$-axis and $t^{\prime}$-axis is the auxiliary origin $O^{\prime}$. If we denote the time coordinate of event $B$ in the coordinate frame $xOt$ as $t_{B}$ and the time coordinate of event $B^{\prime}$ in the coordinate frame $x^{\prime}O^{\prime}t^{\prime}$ as $t^{\prime}_{B^{\prime}}$, then, as a result of such coordinate transformation, $t^{\prime}_{B^{\prime}}=t_{B}$. The method that we evaluate the time coordinate of a real event with the help of an auxiliary event is still valid when dealing with the collision with the moving piston.
@@
Figure 3: Transformation of the coordinate frame and the complete world line. For details, please see the main text of Appendix A.
@@
By performing the coordinate transformation repeatedly, we could figure out the complete auxiliary world line of the particle and every auxiliary coordinate frame. The possible space-time region is $S_{0}=\{(x,t)|0\leq t\leq\tau,0\leq x\leq L+v_{p}t\}$, and the auxiliary regions can be determined accordingly: each gray line segment represents $t_{i}=0$ or $t_{i}=\tau$, $t_{i}$ is the time coordinate of an auxiliary event in the $i$-th auxiliary coordinate frame. In Fig. 2(b), regions between gray line segments are possible auxiliary regions. The red dot $C$ represents the event that a particle happens to collide exactly $i+1$ times when $t=\tau$. The position and time coordinate of C in the initial coordinate frame $xOt$ is denoted as $(x_{C},t_{C})$. Every auxiliary world line $IC$ with the initial condition $I(x_{I},v_{I})$ that passes through $C$ satisfies the critical condition
$x_{I}=x_{C}-v_{I}\cdot t_{C}$ (46)
@@
which is the same as Eq. (30) when $n=i+1$. Such world lines lie between the two orange dashed lines $L_{-}C$ and $L_{+}C$ where $L_{-}$ and $L_{+}$ denote the particles with the initial position $-L$ and $L$. It is clear that the initial positions and velocities of auxiliary world lines satisfy a linear relation if every world line passes through a fixed point $C$, which explains the pictorial intuition mentioned above.
Appendix B Details of the Integration

### v2

@@ II.5 Limit of sast-moving piston
$\begin{split}\left\langle e^{\beta W}\right\rangle&=P_{0}e^{\beta\cdot 0}\\ &+\int_{v_{p}}^{1}\mathrm{d}v\frac{\varphi_{1}(v)e^{-\frac{\beta}{\sqrt{1-v^{2}}}}}{2K_{1}(\beta)(1-v^{2})^{\frac{3}{2}}}e^{\beta W}.\end{split}$ (34)
The first term is approximately equal to 1 ($P(W>0)\ll 1$ and $P_{0}\approx 1$). From the value of $\exp{(-\beta\Delta F)}=1+v_{p}\tau$, one can expect that, although the probability $P(W>0)$ is vanishingly small, the contribution of the second term is nonzero. This result can be demonstrated transparently in the low-temperature limit. Noticing that
$\begin{split}e^{-\beta(1-v^{2})^{-\frac{1}{2}}}e^{\beta W}=&e^{-\beta H_{\lambda(0)}}e^{\beta(H_{\lambda(0)}-H_{\lambda(\tau)})}\\ =&e^{-\beta H_{\lambda(\tau)}}\\ =&e^{-\beta(1-v_{1}^{2})^{-\frac{1}{2}}},\end{split}$ (35)
@@
When $\beta\to\infty$, $e^{-\beta(1-v^{2}_{1}(v))^{-\frac{1}{2}}}\approx e^{-\beta(1+v^{2}_{1}(v)/2)}$, and this term becomes a Gaussian peak, which is a $\delta$ distribution around $v=2v_{p}/(1+v_{p}^{2})$,
$\begin{split}e^{-\beta(1-v^{2}_{1}(v))^{-\frac{1}{2}}}\approx&e^{-\beta}e^{-\frac{\beta}{2}v^{2}_{1}(v)}\\ \approx&e^{-\beta}\sqrt{\frac{2\pi}{\beta}}\delta(v_{1}(v))\\ =&e^{-\beta}\sqrt{\frac{2\pi}{\beta}}\left[1-\left(\frac{2v_{p}}{1+v_{p}^{2}}\right)^{2}\right]\delta\left(v-\frac{2v_{p}}{1+v_{p}^{2}}\right).\end{split}$ (36)
Together with (30) and (19), the second term in Eq. (34) becomes
@@
This result demonstrates that when the piston moves at a very large $v_{p}$, particles with high initial velocities around $2v_{p}/(1+v_{p}^{2})$ contribute most significantly to the exponential work even the probability is extremely small. Note that in Newtonian mechanics particles with initial velocities around $2v_{p}$ contribute most significantly to the exponential work [4, 8], even the probability is extremely small. It can be seen that the results of the relativistic piston model recover those of the nonrelativistic piston model as expected.
III Conclusion
@@
Let us look further into the main results we have obtained. We see that, as a consequence of the relativistic energyvelocity relation, the trapezoid-shaped work distribution no longer has a series of zeros [^1] In fact, in the nonrelativistic regime, $v_{n}(W)$ and $v_{n+1}(W)$ never overlap, that is, there is only one collision number $n$ corresponding to a particular value of $W$. However, with the relativistic energyvelocity relation, a single value $W$ may correspond to different collision numbers.. Moreover, the number of peaks becomes finite because the speed of light places an upper bound on all speeds. The apparent paradox of the fast-moving piston in Ref. [3] can be reformulated as when the piston is moving at the speed of light instead of infinity. No particle would be able to catch up with the piston and the average exponential work is unity whereas the free energy change is nonzero. We would like to point out that although lightlike world lines exist, we cannot make it stop before and after the moving time period, because such a world-line configuration would violate causality. The best we can do is to take the limiting process of letting the speed of the piston approach the speed of light. Then the order of limit becomes crucial, as we have to integrate out the work $W$ to infinity, and then take the speed limit [4]. This limiting procedure ensures the validity of the JE.
@@
In order to observe a distinct deviation of the relativistic work distribution from its nonrelativistic limit, the speed of the piston should be large enough and the temperature of the ideal gas should be extremely high. Taking hydrogen atoms as an example, to observe the features of the relativistic work distribution, the piston should be as fast as about $3\times 10^{6}~\mathrm{m/s}$, and the temperature of the atoms should be about $10^{12}~\mathrm{K}$. When the speed of the piston is about $3\times 10^{5}~\mathrm{m/s}$ (faster than the Parker Solar Probe, which is the fastest object human ever built) and the temperature of the atoms is $10^{11}~\mathrm{K}$ ($10^{4}$ times hotter than the central temperature of the Sun), the deviation of the relativistic work distribution from the nonrelativistic result becomes barely detectable. Even in such a circumstance, the boundary condition is still difficult to realize. Because the energy scale of the kinetic energy of the atoms is much larger than the energy scale of the chemical bond, the boundary cannot be built by any materials we have already discovered. Based on these facts, we conclude that it is difficult to detect the relativistic effects of the work distribution of ideal gas in a piston system with the current experimental techniques.
@@
In summary, we have studied a simple model of a piston and ideal gas in the framework of the special theory of relativity. We obtained an analytical result of the work distribution (18) and verified the JE. Using our result, it was possible to see the deviation of relativistic work distribution (18) from the nonrelativistic one (31) [3]. In principle, these relativistic corrections become non-negligible in the high-temperature and fast-speed limit. However, the range of parameters where relativistic effects are observable would already be far beyond the current experimental techniques. Our results show that the JE holds true in a wide range of systems with generality and serve the pedagogical purpose.
Acknowledgments
@@
This work was supported by the National Natural Science Foundations of China under Grants No. 12375028 and No. 11825501.
Appendix A Covariant description of the trajectory
@@
In Eq. (6), we extended the range of the initial position $x$ from $[0,L]$ to [L,L][-L,L] in order to remove the negative sign related to the direction of $v$. We also notice that in Eq. (21), when $t_{n}$ happens to be $\tau$, the critical initial position and the initial velocity satisfy a linear relation. Here we introduce a coordinate transformation method to describe the trajectory of a particle, which will provide pictorial intuition of the two facts.
@@
To begin with, let us consider the collision between the particle and the fixed boundary. In the reference frame of the fixed boundary (which is the same as the laboratory frame), the coordinate frame consists of an $x$ axis and $t$ axis, and the world line of the moving piston is $l$ (see Fig. 4(a)). We could draw an auxiliary world line of the moving piston, named $l^{\prime}$. The auxiliary world line $l^{\prime}$ and the real world line $l$ are mirror-symmetric about the $t$ axis. The world line of the particle is a polyline II-AA-$B$. Point $I$ denotes the initial condition of the particle, point $A$ denotes the collision event between the particle and the fixed boundary, and point B denotes the next collision event between the particle and the moving piston. Transformation of the particle’s world line works as follows: the auxiliary event of event $B$ is point $B^{\prime}$. Points $B$ and $B^{\prime}$ are mirror symmetric about the $t$ axis. The auxiliary world line of the particle is II-AA-$B^{\prime}$. The real world line’s $AB$ part and the auxiliary world line’s $AB^{\prime}$ part are also mirror symmetric about the $t$ axis. As a result of such transformation, the auxiliary world line of the particle II-AA-$B^{\prime}$ is a straight line, which eliminates the world line’s direction change during a collision. Meanwhile, in the frame of the fixed boundary, the auxiliary collision event $B^{\prime}$ has the same time coordinate as event $B$. Therefore, we could use an auxiliary event to evaluate the time when a real collision takes place.
@@
We may also map the initial condition I(x,v)I(x,-v) to an auxiliary initial condition $I^{\prime}(-x,v)$, which is the same as the treatment for Eq. (6). The real motion is that a particle starts from the initial position $I$ with a positive initial position $x$ but a negative initial velocity v-v (the particle moves away from the moving piston), and then the particle collides with the fixed boundary at event $A$. Meanwhile, the auxiliary motion is that a particle starts from $I^{\prime}$ with a positive initial velocity $v$ but a negative initial position x-x. The auxiliary particle passes through the fixed boundary at event $A$ without any collision. After event $A$, both the particle and the auxiliary particle move towards the moving piston, and finally collide with the piston at the same event $B$. The rest trajectories are equivalent, for both the particle with the initial condition (x,v)(x,-v) and the auxiliary particle with the initial condition (x,v)(-x,v). Because there is no work done during the collision event $A$, it is convenient to extend the range of $x$ from $[0,L]$ to [L,L][-L,L] while limiting the range of $v$ from [c,c][-c,c] to $[0,c]$ in the calculation of the work distribution.
Figure 5: Transformation of a world line. For details see the text Appendix A.
@@
After explaining the treatment in Eq. (6), we now continue to the understanding of Eq. (21). Similar to the mirror symmetric operation in the laboratory reference frame, we could deal with the collision between the particle and the moving piston by carrying out a mirror symmetric operation in the reference frame of the piston. After a Lorentz transformation back to the laboratory reference frame, the result is shown in Fig. 5(a). Points II,AA, and $B$ denote the initial state of the particle, a collision between the particle and the moving piston, and the next collision between the particle and the fixed boundary. Point $B^{\prime}$ is the auxiliary event of $B$. Just like the case in Fig. 4(a), in the reference frame of the moving piston, the auxiliary world line of the particle is a straight line. Therefore, after a Lorentz transformation the auxiliary world line of the particle is still a straight line II-AA-$B^{\prime}$. The auxiliary world line of the fixed boundary is denoted by $t^{\prime}$. The velocity of $t^{\prime}$ is $2v_{p}/(1+v_{p}^{2})$, according to the Lorentz transformation. The intercept of the line $t^{\prime}$ can be determined as follows. The world line of the fixed boundary and moving piston intersect at point $R$ with space and time coordinate $(0,-L/v_{p})$. In the reference frame of the moving piston, the auxiliary world line of the fixed boundary and the moving piston intersect at the same point $R$. Therefore, in the laboratory reference frame, the auxiliary world line of the fixed boundary $t^{\prime}$ also passes through $R$, and the intercept of $t^{\prime}$ must be $L/v_{p}$.
@@
The mirror-symmetric operation transforms not only the world lines, but also the coordinate frames. In the reference frame of the fixed boundary, the $t$ axis is the world line of the boundary itself. After a coordinate transformation related to a collision with the moving piston, the auxiliary world line of the fixed boundary represents the auxiliary $t^{\prime}$ axis. The auxiliary $x^{\prime}$ axis should be orthogonal to the $t^{\prime}$ axis, and the origin of coordinates (point $O$) should be transformed accordingly: The $x$ axis and $l$ intersect at $P_{0}$, and thus the $x^{\prime}$ axis also passes through $P_{0}$ (see Fig. 6(a). The intersection of the auxiliary $x^{\prime}$ axis and the $t^{\prime}$ axis is the auxiliary origin $O^{\prime}$. If we denote the time coordinate of event $B$ in the coordinate frame $xOt$ by $t_{B}$ and the time coordinate of event $B^{\prime}$ in the coordinate frame $x^{\prime}O^{\prime}t^{\prime}$ by $t^{\prime}_{B^{\prime}}$, then, as a result of such coordinate transformation, $t^{\prime}_{B^{\prime}}=t_{B}$. The method that we evaluate the time coordinate of a real event with the help of an auxiliary event is still valid when dealing with the collision with the moving piston.
@@
Figure 6: Transformation of the coordinate frame and the complete world line. For details see the text Appendix A.
@@
By performing the coordinate transformation repeatedly, we could figure out the complete auxiliary world line of the particle and every auxiliary coordinate frame. The possible space-time region is $S_{0}=\{(x,t)|0\leq t\leq\tau,0\leq x\leq L+v_{p}t\}$, and the auxiliary regions can be determined accordingly: Each gray line segment represents $t_{i}=0$ or $t_{i}=\tau$, where $t_{i}$ is the time coordinate of an auxiliary event in the $i$th auxiliary coordinate frame. In Fig. 6(b), regions between gray line segments are possible auxiliary regions. The red dot $C$ represents the event that a particle happens to collide exactly $i+1$ times when $t=\tau$. The position and time coordinate of C in the initial coordinate frame $xOt$ is denoted by $(x_{C},t_{C})$. Every auxiliary world line $IC$ with the initial condition $I(x_{I},v_{I})$ that passes through $C$ satisfies the critical condition
$x_{I}=x_{C}-v_{I}t_{C}$ (38)
@@
which is the same as Eq. (21) when $n=i+1$. Such world lines lie between the two orange dashed lines $L_{-}C$ and $L_{+}C$ where $L_{-}$ and $L_{+}$ denote the particles with the initial positions L-L and $L$, respectively. It is clear that the initial positions and velocities of auxiliary world lines satisfy a linear relation if every world line passes through a fixed point $C$, which explains the pictorial intuition mentioned above.
Appendix B Jarzynski’s equality under the framework of relativistic mechanics
The original proof of the JE was based mainly on the classical Hamilton mechanics and its corollary, i.e., Liouville’s theorem [1]. In this Appendix, we extend the proof to the framework of relativistic mechanics. We take the spacetime dimension to be $1+1$ to suit our model. Note that the generalization to higher spacetime dimensions is straightforward. The manifestly covariant dynamics for a particle of the static mass $m$ can be formulated as
$f^{\mu}=m\frac{\mathrm{d}u^{\mu}}{\mathrm{d}\zeta},$ (39)
$\frac{\partial V}{\partial q}=\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{mu}{\sqrt{1-u^{2}/c^{2}}}\right),$ (40)
where $V$ denotes the potential, $c$ is the speed of light, and $q,u,$ and $t$ are the position, velocity and time, respectively measured in the laboratory frame. With this form of dynamics, we can easily construct the Hamiltonian of a system of $N$ particles of mass $m_{1},\,m_{2},\,...,\,m_{N}$. Furthermore, we let the Hamiltonian be controlled by an external agent via the parameter $\lambda=\lambda(t)$ of the potential $V_{\lambda}.$ The Hamiltonian reads
where $q_{i}$ is the position of the $i$th particle and $p_{i}$ its conjugate momentum.
The canonical equations are
As long as the canonical equations are formulated, one can easily generalize Liouville’s theorem to relativistic regime [9]. The theorem states that the Jacobian determinant of the canonical coordinates $(p_{i,t},q_{i,t})$ at time $t$ as functions of the initial canonical coordinates $(p_{i,0},q_{i,0})$ at time 0 is
The initial equilibrium state of the inverse temperature $\beta$ is determined by a probability distribution $\rho(p_{i,0},q_{i,0})=\exp{[-\beta H(p_{i,0},q_{i,0})]}/Z_{0}$ with $Z_{0}=\int\mathrm{d}p_{i,0}\mathrm{d}q_{i,0}\exp[-\beta H_{\lambda(0)}(p_{i,0},q_{i,0})]$ the initial partition function. We saw in Subsec. II.2) that for ideal gas the distribution is the so-called Maxwell-Jüttner distribution [7]. The proof of the JE follows by calculating the expectation value of the exponential work $\exp{[W(p_{i},q_{i},\tau)]}$ done by the system along the trajectory up to time $\tau.$ By definition we have
where Liouville’s theorem is used for the third equality and $Z_{\tau}$ denotes the final equilibrium state partition function. The partition functions can be expressed in the form of free energy (see, for example, [10]), resulting in
$\left\langle e^{\beta W}\right\rangle=e^{-\beta(F_{\tau}-F_{0})},$ (47)
where $F_{\tau}$ and $F_{0}$ are the free energies of the equilibrium states corresponding to $\lambda(\tau)$ and $\lambda(0)$. Thus, we demonstrate the validity of the JE under the framework of relativistic mechanics.
Appendix C Derivation of Liouville’s Theorem in the Relativistic Piston Model
Here we explain the derivation of Eq. (16) from Eq. (15). The Jacobian determinat is
$\begin{split}\left|\frac{\partial(p_{\tau},x_{\tau})}{\partial(p,x)}\right|&=\left|\frac{\partial x_{\tau}}{\partial x}\frac{\partial p_{\tau}}{\partial p}-\frac{\partial x_{\tau}}{\partial p}\frac{\partial p_{\tau}}{\partial x}\right|\\ &=\left|\frac{\partial x_{\tau}}{\partial x}\frac{\partial p_{\tau}}{\partial p}-\frac{\partial x_{\tau}}{\partial p}0\right|\\ &=(v_{n}+v_{p})\frac{\partial t_{n}}{\partial x}\frac{\mathrm{d}p_{n}}{\mathrm{d}p},\end{split}$ (48)
where
$\begin{split}\frac{\mathrm{d}p_{n}}{\mathrm{d}p}&=\frac{\mathrm{d}p_{n}}{\mathrm{d}v_{n}}\frac{\mathrm{d}v_{n}}{\mathrm{d}v}\left(\frac{\mathrm{d}p}{\mathrm{d}v}\right)^{-1}\\ &=\left(\frac{1-\frac{v^{2}}{c^{2}}}{1-\frac{v_{n}^{2}}{c^{2}}}\right)^{\frac{3}{2}}\frac{4\alpha_{p}^{2n}c^{2}}{(c-v+\alpha_{p}^{2n}(c+v))^{2}}\\ &=\frac{\alpha_{p}^{-n}\left(c-v+\alpha_{p}^{2n}(c+v)\right)}{2c}.\end{split}$ (49)
Noting that $v_{p}=(1-\alpha_{p})c/(1+\alpha_{p}),$ we finally get
$\begin{split}\left|\frac{\partial(p_{\tau},x_{\tau})}{\partial(p,x)}\right|&=\left(v_{n}+\frac{1-\alpha_{p}}{1+\alpha_{p}}c\right)\frac{\partial t_{n}}{\partial x}\\ &\times\frac{\alpha_{p}^{-n}\left(c-v+\alpha_{p}^{2n}(c+v)\right)}{2c}\\ &=\frac{2\left((c+v)\alpha_{p}^{2n}-\alpha_{p}(c-v)\right)}{(\alpha_{p}+1)\left((c+v)\alpha_{p}^{2n}+c-v\right)}\\ &\times\frac{(1+\alpha_{p})\left(\alpha_{p}^{n}-\alpha_{p}^{n+1}\right)c}{(1-\alpha_{p})\left((c+v)\alpha_{p}^{2n}-(c-v)\alpha_{p}\right)}\\ &\times\frac{\alpha_{p}^{-n}\left(c-v+\alpha_{p}^{2n}(c+v)\right)}{2c}\\ &=1,\end{split}$ (50)
which demonstrates that the Jacobian determinant is equal to unity.
Appendix D Details of the Integration